# Measure theoretic properties of large products of consecutive partial quotients

### v1

@@ Abstract.
The theory of uniform approximation of real numbers motivates the study of products of consecutive partial quotients in regular continued fractions. For any non-decreasing positive function $\varphi:\mathbb{N}\to\mathbb{R}^{+}$ and $\ell\in\mathbb{N}$, we determine the Lebesgue measure and Hausdorff dimension of the set $\mathcal{F}_{\ell}(\varphi)$ of irrational numbers $x$ whose regular continued fraction $x\leavevmode\nobreak\ =\leavevmode\nobreak\ [a_{1}(x),a_{2}(x),\ldots]$ is such that for infinitely many $n\in\mathbb{N}$ there are two numbers $1\leq j<k\leq n$ satisfying
$a_{k}(x)\cdots a_{k+\ell-1}(x)\geq\varphi(n),\;a_{j}(x)\cdots a_{j+\ell-1}(x)\geq\varphi(n).$
One of the consequences of the results is that the strong law of large numbers for products of $\ell$ consecutive partial quotients is impossible even if the block with the largest product is removed.
@@ 1. Introduction
Regular continued fractions constitute a representation system of real numbers with outstanding approximation properties. Recall that, for each $x\in\mathbb{R}\setminus\mathbb{Q}$ there is a single sequence of integers $(a_{j}(x))_{j\geq 0}$ with $a_{j}(x)\geq 1$ for $j\geq 1$ such that
@@
The numbers $a_{n}(x)$ are known as the partial quotients of $x$. We have $x\in[0,1)$ if and only if $a_{0}(x)=0$. In this case, we write the continued fractions of $x$ as $[a_{1}(x),a_{2}(x),\ldots]$. It is well known that the $n$-th convergent of $x$ given by
@@
where $q_{n}(x)\geq 1$ and $p_{n}(x)$ are coprime integers, satisfies the relation
@@
see Equation (34) in [16]. Because $(q_{n}(x))_{n\geq 1}$ is strictly increasing, the rate at which rationals approximate $x$ is closely related to the size of its partial quotients. This leads us to consider, for any function $\varphi:\mathbb{N}\to\mathbb{R}_{>0}$, the set
@@
Here and in what follows “i.m.” stands for “infinitely many”. The well-known Borel-Bernstein Theorem establishes the Lebesgue measure $\mathfrak{m}$ of $\mathcal{E}_{1}(\varphi)$:
@@
For proof of this theorem, see for example [2, Theorem 1.11]. Łuczak [21] and Feng et al. [6] calculated the Hausdorff dimension, denoted as $\dim_{\mathrm{H}}$, of $\mathcal{E}_{1}(\varphi)$ when $\varphi(n)=c^{b^{n}}$ for some fixed $c,b>1$ (see Lemma 2.7). Wang and Wu solved the general case in [27] (see Theorem 1.2 below with $\ell=1$).
@@
From a probabilistic perspective, we may interpret partial quotients as random variables defined on $[0,1)$ (their definition of rational numbers is irrelevant for our purposes). These random variables are not independent, hence the Borel-Bernstein Theorem is not a trivial consequence of the Borel-Cantelli lemma. Moreover, the Borel-Bernstein theorem tells us that for almost every number $x$ (with respect to $\mathfrak{m}$) the inequality $a_{n}(x)\geq n\log n$ holds infinitely often. This implies that, writing $S_{n}(x):=a_{1}(x)+\cdots+a_{n}(x)$, we have
@@
Despite not being independent and having infinite expected value, the partial quotients satisfy a weak law of large numbers. Khinchin [17, §. 4] proved that $S_{n}/(n\log n)$ converges in Lebesgue measure to $\frac{1}{\log 2}$ as $n\to\infty$. Furthermore, Philipp [23, Theorem 1] proved that there is no reasonable function $\sigma$ for which $S_{n}/\sigma(n)$ converges almost everywhere to a finite positive constant. Large partial quotients are to blame for this phenomenon. Certainly, Diamond and Vaaler [4, Corollary 1] showed that convergence almost everywhere does hold for Khinchin’s averaging function when omitting the largest partial quotient: for almost every $x\in[0,1)$ we have
@@
This result was extended by Hu, Hussain, and Yu for sums of products of two partial quotients [10, Theorem 1.5]. They showed that almost every $x\in[0,1)$ satisfies
@@
Recently, Tan, Tian, and Wang [25] generalised this result of Diamond and Vaaler to a wider family of functions. Later, Tan and Zhou [26] also extended the result of Hu, Hussain and Yu. More precisely, define
@@
Our first result, which includes Theorem 1.4, completes the picture of the Lebesgue measure of $\mathcal{F}_{\ell}(\varphi)$.
@@
Let $\varphi:\mathbb{N}\to[2,\infty)$ be non-decreasing. For $\ell\geq 2$, the Lebesgue measure of $\mathcal{F}_{\ell}(\varphi)$ is given by
$\mathfrak{m}(\mathcal{F}_{\ell}(\varphi))=\begin{cases}0&\text{\rm if }\displaystyle\sum_{n\geq 1}\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}<\infty,\\ 1&\text{\rm if }\displaystyle\sum_{n\geq 1}\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}=\infty.\end{cases}$
@@
Since all the partial quotients are at least $1$, for almost every $x\in[0,1)$ we have
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}a_{j}(x)\cdots a_{j+\ell}(x)=\infty.$
Theorem 1.5 implies that for almost every $x\in[0,1)$ there are infinitely many numbers $n\in\mathbb{N}$ for which there are $1\leq j<k\leq n$ satisfying
$a_{j}(x)\cdots a_{j+\ell-1}(x)\geq n\log^{\ell-1}n\;\text{ and }\;a_{k}(x)\cdots a_{k+\ell-1}(x)\geq n\log^{\ell-1}n.$
@@
$\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{j=1}^{n}a_{j}(x)\cdots a_{j+\ell-1}(x)-\max_{1\leq j\leq n}a_{j}(x)\cdots a_{j+\ell-1}(x)\right)=\infty.$
@@
$\lim_{n\to\infty}\frac{1}{n\log^{\ell}n}\left(\sum_{j=1}^{n}a_{j}(x)\cdots a_{j+\ell}(x)-\max_{1\leq j\leq n}a_{j}(x)\cdots a_{j+\ell-1}(x)\right)=\frac{1}{\ell\log 2}.$
@@
While proving Theorem 1.5 we must obtain new estimates corresponding to overlapping blocks of consecutive partial quotients with large products. The number of terms overlapping in Theorem 1.5 ranges from 00 up to $\ell-1$ whereas in Theorem 1.4 the only overlaps are of length 00 and $1$ and there are no overlaps in Theorem 1.3.
@@
The paper is organised as follows. Section 2 contains some preliminaries and auxiliary lemmas from measure theory and continued fractions, in particular on the Hausdorff dimension and pressure functions. In Section 3 we estimate certain series that are fundamental for our proofs. The proof of Theorem 1.5 is in Section 4 and the proof of Theorem 1.10 is in Section 5. Lastly, Section 6 contains some final remarks and results developed to further this study.
@@
Acknowledgements. GGR and MH were funded by the Australian Research Council discovery project grant number 200100994. We thank Nikita Shulga and Ben Ward for useful discussions.
@@ 2.1. Full and null sets
The Borel-Cantelli Lemma is a fundamental tool for determining the measure of a limsup set. A proof may be found, for example, in [15, Lemma 4.18].
@@
The independence assumption is rather restrictive, it will not hold in our context. A way to circumvent this problem is proving the positive measure of the limsup set and showing by some other means that the measure is either full or null.
We show the full measure statement in Theorem 1.5 using a ubiquity-type property and Lemma 2.2. This lemma is attributed to Knopp [19] by Tan and Zhou [26]. Recall that, given two measurable sets $A,B$ in a probability space $(X,\mathscr{A},\mu)$, we write $A\equiv B\pmod{\mu}$ if and only if $\mu(A\setminus B)=\mu(B\setminus A)=0$.
@@
Let $A\subset[0,1)$ be a Borel set and $\mathcal{C}$ a class of subintervals of $[0,1)$ satisfying both of the following conditions:
@@
    There is a constant $c>0$ such that for any $B\in\mathcal{C}$ we have $\mathfrak{m}(A\bigcap B)\geq c\mathfrak{m}(B)$.
@@
In order to apply Lemma 2.2, we need the following result due to Chung and Erdös [3] (see also [24, Lemma 6] for the formulation stated below).
Lemma 2.3 (Chung, Erdös, [3]).
@@ 2.2. Hausdorff dimension
Given a non-empty subset $A$ of real numbers, let $|A|$ be its diameter. For $\delta>0$, by a $\delta$-cover of $A$ we mean a countable collection $\{B_{i}:i\in\mathbb{N}\}$ such that
@@ 2.3. Continued fractions
An elementary result (see, for example, [9, Proposition 1.1]) states that the sequences $(p_{n}(x))_{n\geq 0}$, $(q_{n}(x))_{n\geq 1}$ verify
@@ 2.4. Pressure functions
The arguments of [27, Section 2] allow us to conclude
@@
Moreover, relying on Theorem 1.8, [25, Proposition 2.11], and $g_{3}(x)\leq 3x-1$ for $x\geq\frac{2}{5}$ we may show that $p_{B}\geq\frac{1}{2}$.
The lower estimate for the Hausdorff dimension of sets similar to that in Theorem 1.10 tends to be computationally demanding. A general strategy is to construct a Cantor set $C$ contained in our set of interest, afterwards define a probablity measure on $C$ and apply the Mass Distribution Principle [5, Proposition 4.2]. Recently, Hussain and Shulga [14] computed the Hausdorff dimension of a large family of sets defined by continued fractions restrictions. For $A_{0},\ldots,A_{\ell-1}>1$, define
@@ 3. Series estimates
Let us start with the lemmas needed in the proof of Theorem 1.5. The first one can be retrieved from the proof of [11, Theorem 1.5].
@@
We recall a result from analytic number theory (see, for example, [8, Section 4]).
@@
(6) $\sum_{\begin{subarray}{c}a_{1}\cdots a_{r}b_{1}\cdots b_{j}\geq M\\ b_{1}\cdots b_{j}c_{1}\cdots c_{r}\geq M\end{subarray}}\frac{1}{a_{1}^{2}\cdots a_{r}^{2}b_{1}^{2}\cdots b_{j}^{2}c_{1}^{2}\cdots c_{r}^{2}}\asymp_{\ell}\frac{\log^{j-1}(M)}{M},$
@@
$\asymp\sum_{b_{1}\cdots b_{j}\leq\frac{M}{2}}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\left(\frac{1}{\frac{M}{b_{1}\cdots b_{j}}}\right)^{2}$
$=\frac{1}{M^{2}}\sum_{b_{1}\cdots b_{j}\leq\frac{M}{2}}1\asymp\frac{\log^{j-1}M}{M}.$
@@
$\sum_{\frac{M}{2}\leq a_{2}\cdots a_{\ell}\leq M}\sum_{a_{1}\leq\frac{M}{a_{2}\ldots a_{\ell}}}\sum_{a_{\ell+2}\geq\frac{M}{a_{1}\cdots a_{\ell}}}\frac{1}{a_{1}^{2}\cdots a_{\ell+1}^{2}}$ $\asymp\sum_{a_{2}\cdots a_{\ell}\leq M}\frac{1}{a_{2}^{2}\cdots a_{\ell}^{2}}$
$\asymp\sum_{a_{2}\cdots a_{\ell}\leq\frac{M}{2}}\frac{1}{a_{2}^{2}\cdots a_{\ell}^{2}}\asymp\frac{\log^{\ell-2}M}{M}.$
@@
For any $n\in\mathbb{N}_{\geq\ell}$ and $k\in\{1,\ldots,n-1\}$, consider

### v2

@@ Abstract.
The theory of uniform approximation of real numbers motivates the study of products of consecutive partial quotients in regular continued fractions. For any non-decreasing positive function $\varphi:\mathbb{N}\to[2,\infty)$, we determine the Lebesgue measure and Hausdorff dimension of the set $\mathcal{F}_{\ell}(\varphi)$ of irrational numbers $x$ whose regular continued fraction $x\leavevmode\nobreak\ =\leavevmode\nobreak\ [a_{1}(x),a_{2}(x),\ldots]$ is such that for infinitely many $n\in\mathbb{N}$ there are two numbers $1\leq j<k\leq n$ satisfying
$a_{k}(x)a_{k+1}(x)a_{k+2}(x)\geq\varphi(n),\;a_{j}(x)a_{j+1}(x)a_{j+2}(x)\geq\varphi(n).$
One of the consequences of the results is that the strong law of large numbers for products of $3$ consecutive partial quotients is impossible even if the block with the largest product is removed.
@@ 1. Introduction
Regular continued fractions are a representation system of real numbers with outstanding approximation properties. Recall that, for each $x\in\mathbb{R}\setminus\mathbb{Q}$ there is a single sequence of integers $(a_{j}(x))_{j\geq 0}$ with $a_{j}(x)\geq 1$ for $j\geq 1$ such that
@@
The numbers $a_{n}(x)$ are known as the partial quotients of $x$. We have $x\in[0,1)$ if and only if $a_{0}(x)=0$. In this case, we write the continued fractions of $x$ as $[a_{1}(x),a_{2}(x),\ldots]$. It is well known that the $n$-th convergent of $x$, given by
@@
where $q_{n}(x)\geq 1$ and $p_{n}(x)$ are coprime integers, satisfies
@@
(see, for example, [17, Equation (34)]). Because $(q_{n}(x))_{n\geq 1}$ is strictly increasing, the rate at which rationals approximate $x$ is closely related to the size of its partial quotients. This leads us to consider, for any function $\varphi:\mathbb{N}\to\mathbb{R}_{>0}$, the set
@@
Here and in what follows, “i.m.” stands for “infinitely many”. The well-known Borel-Bernstein Theorem establishes the Lebesgue measure $\mathfrak{m}$ of $\mathcal{E}_{1}(\varphi)$:
@@
For proof of this theorem, see, for example, [2, Theorem 1.11]. Łuczak [22] and Feng et al. [7] calculated the Hausdorff dimension, denoted as $\dim_{\mathrm{H}}$, of $\mathcal{E}_{1}(\varphi)$ when $\varphi(n)=c^{b^{n}}$ for some fixed $c,b>1$ (see Lemma 2.7). Wang and Wu solved the general case in [29] (see Theorem 1.2 below with $\ell=1$).
@@
From a probabilistic perspective, we may interpret partial quotients as random variables defined on $[0,1)$ (their definition of rational numbers is irrelevant for our purposes). These random variables are not independent, hence the Borel-Bernstein Theorem is not a trivial consequence of the Borel-Cantelli lemma. Moreover, the Borel-Bernstein theorem tells us that the inequality $a_{n}(x)\geq n\log n$ holds infinitely often for almost every number $x$ (with respect to $\mathfrak{m}$). This implies that, writing $S_{n}(x):=a_{1}(x)+\cdots+a_{n}(x)$, we have
@@
Despite not being independent and having infinite expected value, the partial quotients satisfy a weak law of large numbers. Khinchin [18, §. 4] proved that $S_{n}/(n\log n)$ converges in Lebesgue measure to $\frac{1}{\log 2}$ as $n\to\infty$. Furthermore, Philipp [24, Theorem 1] proved that there is no reasonable function $\sigma$ for which $S_{n}/\sigma(n)$ converges almost everywhere to a finite positive constant. Large partial quotients are to blame for this phenomenon. Certainly, Diamond and Vaaler [4, Corollary 1] showed that convergence almost everywhere does hold for Khinchin’s averaging function when omitting the largest partial quotient: for almost every $x\in[0,1)$ we have
@@
This result was extended by Hu, Hussain, and Yu for sums of products of two partial quotients [11, Theorem 1.5]. They showed that almost every $x\in[0,1)$ satisfies
@@
Recently, Tan, Tian, and Wang [27] generalized this result of Diamond and Vaaler to a wider family of functions. Later, Tan and Zhou [28] also extended the result of Hu, Hussain and Yu. More precisely, define
@@
Our first result extends Theorem 1.4 to blocks of length $3$.
@@
Let $\varphi:\mathbb{N}\to[2,\infty)$ be non-decreasing. The Lebesgue measure of $\mathcal{F}_{3}(\varphi)$ is given by
$\mathfrak{m}(\mathcal{F}_{3}(\varphi))=\begin{cases}0&\text{\rm if }\displaystyle\sum_{n\geq 1}\frac{n\log^{4}\varphi(n)}{\varphi^{2}(n)}+\frac{\log\varphi(n)}{\varphi(n)}<\infty,\\ 1&\text{\rm if }\displaystyle\sum_{n\geq 1}\frac{n\log^{4}\varphi(n)}{\varphi^{2}(n)}+\frac{\log\varphi(n)}{\varphi(n)}=\infty.\end{cases}$
@@
Using Birkhoff’s Ergodic Theorem, we may show that for almost every $x\in[0,1)$ we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}a_{j}(x)\cdots a_{j+\ell}(x)=\infty$
(see, for example, [5, Corollary 3.8]). Moreover, Theorem 1.5 implies that for almost every $x\in[0,1)$ there are infinitely many numbers $n\in\mathbb{N}$ for which there are $1\leq j<k\leq n$ satisfying
$a_{j}(x)a_{j+1}(x)a_{j+2}(x)\geq n\log^{2}n\;\text{ and }\;a_{k}(x)a_{k+1}(x)a_{k+2}(x)\geq n\log^{2}n.$
@@
$\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{j=1}^{n}a_{j}(x)a_{j+1}(x)a_{j+2}(x)-\max_{1\leq k\leq n}a_{k}(x)a_{k+1}(x)a_{k+2}(x)\right)=\infty.$
@@
$\lim_{n\to\infty}\frac{1}{n\log^{3}n}\left(\sum_{j=1}^{n}a_{j}(x)a_{j+1}(x)a_{j+2}(x)-\max_{1\leq k\leq n}a_{k}(x)a_{k+1}(x)a_{k+2}(x)\right)=\frac{1}{3\log 2}.$
@@
While proving Theorem 1.5 we must obtain new estimates corresponding to overlapping blocks of consecutive partial quotients with large products. The number of terms overlapping in Theorem 1.5 ranges from 00 up to $\ell-1=2$ whereas in Theorem 1.4 the only overlaps are of length 00 and $1$ and there are no overlaps in Theorem 1.3.
@@
The paper is organized as follows. Section 2 contains some preliminaries and auxiliary lemmas from measure theory and continued fractions, in particular on the Hausdorff dimension and pressure functions. In Section 3 we estimate certain series that are fundamental for our proofs. The proof of Theorem 1.5 is in Section 4 and the proof of Theorem 1.10 is in Section 5. Lastly, Section 6 contains some final remarks and results developed to further this study.
@@
Acknowledgements. GGR and MH were funded by the Australian Research Council discovery project grant number 200100994. We thank Nikita Shulga, Valeriia Starichkova, and Ben Ward for useful discussions.
@@ 2.1. Full and null sets
The Borel-Cantelli Lemma is a fundamental tool for determining the measure of limsup sets. A proof may be found, for example, in [16, Lemma 4.18].
@@
The independence assumption is rather restrictive, it will not hold in our context. A way to circumvent this problem is proving the positive measure of the limsup set and showing by some other means that the measure is either full or null. We show the full measure statement in Theorem 1.5 using a ubiquity-type property and Lemma 2.2. This lemma is attributed to Knopp [20] by Tan and Zhou [28].
Given two measurable sets $A,B$ in a probability space $(X,\mathscr{A},\mu)$, we write $A\equiv B\pmod{\mu}$ if and only if $\mu(A\setminus B)=\mu(B\setminus A)=0$.
@@
Let $A\subseteq[0,1)$ be a Borel set and $\mathcal{C}$ a class of subintervals of $[0,1)$ satisfying both of the following conditions:
@@
    There is a constant $c>0$ such that for any $B\in\mathcal{C}$ we have $\mathfrak{m}(A\cap B)\geq c\mathfrak{m}(B)$.
@@
In order to apply Lemma 2.2, we need the following result due to Chung and Erdős [3] (see also [26, Lemma 6] for the formulation stated below).
Lemma 2.3 (Chung, Erdős, [3]).
@@ 2.2. Hausdorff dimension
Given a non-empty subset $A$ of real numbers, let $|A|$ be its diameter. For $\delta>0$, by a $\delta$-cover of $A$ we mean a countable collection of subsets of real numbers $\{B_{i}:i\in\mathbb{N}\}$ such that
@@ 2.3. Continued fractions
An elementary result (see, for example, [10, Proposition 1.1]) states that the sequences $(p_{n}(x))_{n\geq 0}$, $(q_{n}(x))_{n\geq 1}$ verify
@@ 2.4. Pressure functions
The arguments of [29, Section 2] allow us to conclude
@@
Moreover, relying on Theorem 1.8, [27, Proposition 2.11], and $g_{3}(x)\leq 3x-1$ for $x\geq\frac{2}{5}$ we may show that $p_{B}\geq\frac{1}{2}$.
The lower estimate for the Hausdorff dimension of sets similar to that in Theorem 1.10 tends to be computationally demanding. A general strategy is to construct a Cantor set $C$ contained in our set of interest, afterwards define a probablity measure on $C$ and apply the Mass Distribution Principle [6, Proposition 4.2]. Recently, Hussain and Shulga [14] computed the Hausdorff dimension of a large family of sets defined by continued fractions restrictions. For $A_{0},\ldots,A_{\ell-1}>1$, define
@@ 3. Series estimates
Let us start with the lemmas needed in the proof of Theorem 1.5. The first one can be retrieved from the proof of [12, Theorem 1.5].
@@
We recall a result from analytic number theory (see, for example, [9, Section 4]).
@@
For any $j\in\mathbb{N}$ and large $M>1$, we have
$\sum_{\begin{subarray}{c}ab_{1}b_{2}\cdots b_{j}\geq M\\ b_{1}b_{2}\cdots b_{j}c\geq M\end{subarray}}\frac{1}{a^{2}b_{1}^{2}\cdots b_{j}^{2}c^{2}}\asymp_{j}\frac{\log^{j-1}M}{M},$
Proof.
Note that
$\sum_{\begin{subarray}{c}ab_{1}\cdots b_{j}\geq M\\ b_{1}\cdots b_{j}c\geq M\end{subarray}}\frac{1}{a^{2}b_{1}^{2}\cdots b_{j}^{2}c^{2}}$ $=\sum_{b_{1}\cdots b_{j}\geq M}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\left(\sum_{a>\frac{M}{b_{1}\cdots b_{j}}}\frac{1}{a^{2}}\right)^{2}+\sum_{b_{1}\cdots b_{j}<M}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\left(\sum_{a>\frac{M}{b_{1}\cdots b_{j}}}\frac{1}{a^{2}}\right)^{2}$
When $b_{1}\cdots b_{j}\geq M$, the restriction $a\geq\frac{M}{b_{1}\cdots b_{j}}$ is irrelevant, so
$\sum_{b_{1}\cdots b_{j}\geq M}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\left(\sum_{a>\frac{M}{b_{1}\cdots b_{j}}}\frac{1}{a^{2}}\right)^{2}\asymp\sum_{b_{1}\cdots b_{j}\geq M}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\asymp\frac{\log^{j-1}M}{M}.$
For the second term, we use the Dirichlet-Piltz formula to obtain
$\sum_{b_{1}\cdots b_{j}<M}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\left(\sum_{a>\frac{M}{b_{1}\cdots b_{j}}}\frac{1}{a^{2}}\right)^{2}\ll\sum_{b_{1}\cdots b_{j}<M}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\frac{1}{\frac{M^{2}}{b_{1}^{2}\cdots b_{j}^{2}}}=\frac{1}{M^{2}}\sum_{b_{1}\cdots b_{j}<M}1\asymp\frac{\log^{j-1}M}{M}.$
Lemma 3.4.
For large $M>1$, we have
(6) $\sum_{\begin{subarray}{c}a_{1}a_{2}b\geq M\\ bc_{1}c_{2}\geq M\end{subarray}}\frac{1}{a_{1}^{2}a_{2}^{2}b^{2}c_{1}^{2}c_{2}^{2}}\asymp\frac{1}{M}.$
Proof.
Observe that
$\sum_{\begin{subarray}{c}a_{1}a_{2}b\geq M\\ bc_{1}c_{2}\geq M\end{subarray}}\frac{1}{a_{1}^{2}a_{2}^{2}b^{2}c_{1}^{2}c_{2}^{2}}=\sum_{b\geq M}\frac{1}{b^{2}}\left(\sum_{a_{1}a_{2}\geq\frac{M}{b}}\frac{1}{a_{1}^{2}a_{2}^{2}}\right)^{2}+\sum_{b<M}\frac{1}{b^{2}}\left(\sum_{a_{1}a_{2}\geq\frac{M}{b}}\frac{1}{a_{1}^{2}a_{2}^{2}}\right)^{2}.$
The condition $a_{1}a_{2}\geq\frac{M}{b}$ is redundant when $b\geq M$, so
$\sum_{b\geq M}\frac{1}{b^{2}}\left(\sum_{a_{1}a_{2}\geq\frac{M}{b}}\frac{1}{a_{1}^{2}a_{2}^{2}}\right)^{2}\asymp\sum_{b\geq M}\frac{1}{b^{2}}\asymp\frac{1}{M}.$
For the second term, we have
$\sum_{b<M}\frac{1}{b^{2}}\left(\sum_{a_{1}a_{2}\geq\frac{M}{b}}\frac{1}{a_{1}^{2}a_{2}^{2}}\right)^{2}\asymp\sum_{b<M}\frac{1}{b^{2}}\frac{\log^{2}\frac{M}{b}}{\frac{M^{2}}{b^{2}}}=\frac{1}{M^{2}}\sum_{b<M}\log^{2}\left(\frac{M}{b}\right)\asymp\frac{1}{M^{2}}\int_{1}^{M}\log^{2}\left(\frac{x}{M}\right)\asymp\frac{1}{M}.$
Lemma 3.5.
@@
(7) $\sum_{\begin{subarray}{c}a_{1}\cdots a_{r}b_{1}\cdots b_{j}\geq M\\ b_{1}\cdots b_{j}c_{1}\cdots c_{r}\geq M\end{subarray}}\frac{1}{a_{1}^{2}\cdots a_{r}^{2}b_{1}^{2}\cdots b_{j}^{2}c_{1}^{2}\cdots c_{r}^{2}}\ll_{\ell}\frac{\log^{2(j-1)}(M)}{M},$
@@
$\asymp\sum_{b_{1}\cdots b_{j}\leq\frac{M}{2}}\frac{1}{b_{1}^{2}\cdots b_{j}^{2}}\left(\frac{\log^{j-1}\left(\frac{M}{b_{1}\cdots b_{j}}\right)}{\frac{M}{b_{1}\cdots b_{j}}}\right)^{2}$
$\asymp\frac{1}{M^{2}}\sum_{b_{1}\cdots b_{j}\leq\frac{M}{2}}\log^{2(j-1)}\left(\frac{M}{b_{1}\cdots b_{j}}\right)$
$\ll\frac{\log^{2(j-1)}M}{M^{2}}\sum_{b_{1}\cdots b_{j}\leq\frac{M}{2}}1$
$\ll\frac{\log^{2(j-1)}M}{M}.$
@@
$\sum_{\frac{M}{2}\leq a_{2}\cdots a_{\ell}\leq M}\sum_{a_{1}\leq\frac{M}{a_{2}\ldots a_{\ell}}}\sum_{a_{\ell+2}\geq\frac{M}{a_{1}\cdots a_{\ell}}}\frac{1}{a_{1}^{2}\cdots a_{\ell+1}^{2}}$ $\asymp\sum_{\frac{M}{2}\leq a_{2}\cdots a_{\ell}\leq M}\frac{1}{a_{2}^{2}\cdots a_{\ell}^{2}}$
$\ll\sum_{a_{2}\cdots a_{\ell}\leq\frac{M}{2}}\frac{1}{a_{2}^{2}\cdots a_{\ell}^{2}}\asymp\frac{\log^{\ell-2}M}{M}.$
@@
We show a slightly different statement for the convergence part which yields the convergence part in Theorem 1.4. To this end, for $\ell\in\mathbb{N}_{\geq 3}$, $n\in\mathbb{N}_{\geq\ell}$, and $k\in\{1,\ldots,n-1\}$, define